OpenStudy (anonymous):

On a scale drawing, 1 cm represents 4 meters. What length on the drawing would be used to represent 6.5 meters?

OpenStudy (anonymous):

1x=4 ax=6.5 a*4=6.5 a=6.5/4=13/8=1&5/8cm (or 1.625cm)

OpenStudy (anonymous):

I don't get it what does it equal to?

OpenStudy (anonymous):

a is the distance in cm that equals 6.5m in the scale drawing

OpenStudy (anonymous):

So it is equal to 1.625 cm?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!